how much time does stone take to reach on earth surface drops from 20m height and what will be the acceleration due to gravity after 3sec?


1
Expert's answer
2020-11-16T06:36:08-0500
"h=0.5gt^2\\20=0.5(9.8)t^2\\t=2.0\ s\\g=9.8\frac{m}{s^2}"